# Iterative Data-Consistent Inversion with Multiple Push-forward Constraints

### v1

@@ 1 Introduction
A key objective of modern computational science and engineering is the rigorous quantification of uncertainty in physics-based models, driven by the desire to produce data-informed predictions. Central to this effort is the formulation and solution of stochastic inverse problems (SIPs), which seek to constrain input uncertainties using observational data. The SIP is broadly defined in the context of this work as constructing a probability measure on an input space ($\Lambda$) of a measurable quantities of interest (QoI) map ($\phi$) that is subject to a push-forward constraint associated with some specified observed probability measure ($P_{\text{obs}}$) on the output space ($\mathcal{D}$). More formally, the solution is a probability measure $P$ on $\Lambda$ such that $P\circ\phi^{-1}=P_{\text{obs}}$. In other words, the solution is a pullback measure of the observed distribution, assigning the correct probability to events in the parameter space based on observed frequencies in the data space.
While the standard SIP considers a single (possibly vector-valued) QoI map, many practical applications involve conditioning parameter uncertainties on multiple, distinct observational datasets associated with separate QoI maps. Such situations arise when different experimental campaigns lead to asynchronous data for which no joint structures are observed or known simultaneously across every QoI component from each QoI map. This leads to the Generalized Stochastic Inverse Problem (GSIP), which requires finding a single probability measure $P$ that simultaneously satisfies $k$ distinct push-forward constraints, $P\circ\phi_{i}^{-1}=P_{\text{obs},i}$, for $i=1,\dots,k$, where $\phi_{i}$ denotes the $i$th (possibly vector-valued) QoI map and $P_{\text{obs},i}$ is the associated observed probability measure. The GSIP presents unique challenges compared to the SIP, particularly because the full joint observed distribution across all output spaces ($\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{k}$ where $\mathcal{D}_{i}$ is the output space associated with $\phi_{i}$) is generally not assumed to be available or necessary for solution. Subsequently, solving the GSIP requires an approach that resolves these constraints individually and iteratively.
@@ 2 Data-Consistent Inversion (DCI): Background, Notation, and Terminology
Let $P_{\text{init}}$ denote a probability measure on $(\Lambda,\mathcal{B}_{\Lambda})$ and $P_{\text{pred}}:=P_{\text{init}}\circ\phi^{-1}$ denote its pushforward measure on $(\mathcal{D},\mathcal{B}_{\mathcal{D}})$, then the disintegration of $P_{\text{init}}$ yields
@@ 3 DCI Solutions Minimize f-Divergences
It is worth noting that by varying $P_{\text{init}}$, we can generate a range of possible solutions, $P_{\text{up}}$, implying that any pullback measure can be recovered by a well chosen initial distribution. It may also be necessary to vary $P_{\text{init}}$ to ensure that $P_{\text{init}}\ll P_{\text{up}}$, which ensures that $\tilde{\mathbb{P}}$ is nonempty. This flexibility in the choice of an initial distribution ultimately allows for a more comprehensive exploration of the solution space. Although varying $P_{\text{init}}$ is not the focus of this work, it will be the topic of a future work as mentioned in Section 7.
Once $P_{\text{init}}$ is chosen, we prove below that the DCI solution is optimal in the sense of minimizing the f-divergences over the space of all probability measures that are absolutely continuous with respect to $P_{\text{init}}$ and whose push-forward measure is equal to $P_{\text{obs}}$. The f-divergences [33] are commonly employed in probability and information theory to quantify the difference between probability measures often in the context of determining optimal parameters or hyper-parameters of density models, e.g., see [25, 5]. A key feature of f-divergences are their flexibility, as they generalize and connect many well-known divergence measures, such as the Kullback-Leibler divergence and total variation distance [27, 41, 5, 1]. Utilizing the notation of this work, the f-divergence is defined as:
@@
where $P\ll P_{\text{init}}$ and $f:[0,\infty)\to\mathbb{R}$ is a convex function satisfying $f(1)=0$ and $f(0)=\lim_{t\rightarrow 0^{+}}f(t)$.
@@
$P(\phi^{-1}(B))=P_{\text{obs}}(B)=\int_{B}\,dP_{\text{obs}}(\omega)=\int_{B}\cfrac{dP_{\text{obs}}}{dP_{\text{pred}}}(\lambda)dP_{\text{pred}}(\omega).$
@@ 4 The Generalized Stochastic Inverse Problem
It is self-evident that if each $P_{\text{obs},i}$ is the induced push-forward measure of some probability measure defined on $(\Lambda,\mathcal{B}_{\Lambda})$, then a solution is guaranteed to exist. We make such an assumption in this work and refer to this measure as the data-generating distribution (denoted as $P_{\text{DG}}$), given its role in inducing the observed probability measures.
We briefly emphasize several important differences between the GSIP and the SIP. First, for the SIP solved by DCI, only a single (vector-valued) map $\phi$ is considered. By contrast, there are multiple (vector-valued) maps considered in the GSIP. Moreover, for the GSIP, it is possible that some components of a map $\phi_{i}$ may also appear in other maps $\phi_{j}$. This is related to a second key difference in terms of how uncertainties in output data are assumed to be quantified in the SIP and GSIP. In the SIP, a single joint observed measure is assumed to be defined on the full data space defined by all the outputs of the map $\phi$. However, the GSIP does not assume that the full joint observed distribution is available on $\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{k}$. A final important note is that a predictability assumption is sufficient to guarantee the DCI solution exists to the SIP (e.g., that $P_{\text{obs}}$ is absolutely continuous with respect to $P_{\text{pred}}=P_{\text{init}}\circ\phi^{-1}$). While predictability assumptions are also required for the solution of the GSIP, a further assumption is needed to guarantee that each of the separate $k$ push-forward constraints of (14) are simultaneously satisfied, which is handled in this work by the additional assumption of a data-generating distribution.
@@ 5 An Iterative DCI Approach
Before we present the proof, we provide high-level comments to help provide some intuition and orient the reader on the steps involved. The proof we provide shows the $k=3$ case to make it clear how these prior convergence results, and we note that the approach we provide is extended to the $k>3$ cases naturally. The first step in the proof constructs a product data space from the $k$ push-forward constraints along with a family of push-forward probability measures since each constraint defines a known marginal on this space. The second step in the proof involves expressing the Radon-Nikodym derivatives of the push-forward measures with respect to the push-forward of $P^{0}=P_{\text{init}}$, which provides a representation in the form of an IPFP on the product data space. To help build some intuition about the structure of these Radon-Nikodym derivatives and the relation to the IPFP method, suppose we equip $\mathbb{R}^{3}$ with the initial density $f^{0}(x,y,z)$ but that we wish to adjust this joint density so that the $x$-marginal is changed from $f^{0}_{1}(x):=\int_{\mathbb{R}^{2}}f^{0}(x,y,z)\,dy\,dz$ to $f_{1}(x)$. This is easily accomplished through the following formula (assuming the support of $f_{1}(x)$ is a subset of the support of $f^{0}_{1}(x)$)
@@ 6.1.1 Numerical Setup
For the algorithm, the stopping criteria were defined as follows. The maximum of number of epochs was set to 100. The tolerance was set to 0.1 for the numerical diagnostic, i.e., if the sample average of the $r$-values deviates by more than 10% from the theoretically expected value of unity, then the algorithm terminates. The tolerance for the KL divergence of all QoI marginals associated with the observed distribution and push-forward of the updated distribution after each epoch was set to ${1}\mathrm{E}{-7}$. In other words, this tolerance is reached at epoch $k$ only if every QoI marginal associated with the observed distribution and the push-forward of the updated distribution is less than ${1}\mathrm{E}{-7}$. Finally, the relative tolerance for the difference in such KL divergences obtained between successive epochs was set to ${1}\mathrm{E}{-2}$. In other words, this tolerance is reached at epoch $k+1$ only if the KL divergences of every QoI marginal associated with the observed and updated distributions are within $1\%$ of those computed at epoch $k$.
@@ 6.1.2 Iterative Results
The relative tolerance stopping criteria was triggered at the 46th epoch where the push-forward of the updated distribution at this epoch produced marginal QoI densities that both had KL divergences from the associated observed marginals that were within 1% from the values computed from the 45th epoch. We illustrate some of the results in Figures 3-5 below. For reference, the top-right plot in each of these figures shows the KDE of the data-generating samples shown as the orange pluses in the left plot of Figure 2. The top-left and top-middle plots in each of these figures show the weighted KDE computed on the initial samples shown as the blue circles in Figure 2 with weights determined at the first (left) and second (middle) iteration of each epoch. The bottom two plots in each of these figures show the KDE estimates of the marginal densities on $Q_{1}$ (left) and $Q_{2}$ (right) associated with the observed distribution and push-forward of the updated distribution at the end of each epoch.
@@
Numerical evidence of convergence is further supported by Figure 6, which shows the rapid decrease in the KL divergence between the QoI marginals associated with the observed distribution and the push-forward of the updated distribution at each iteration. Here, we show results for each iteration (of which there are 92 total) instead of each epoch (of which there are 46) to highlight the impact of alternating between the $Q_{1}$ and $Q_{2}$ distributions at each iteration within an epoch that produce oscillations in the decay trend. The magnitude of the oscillations is observed to quickly decay and is virtually undetectable, at the scale plotted, after 30-40 iterations (i.e., after 15-20 epochs).
@@ 6.1.4 Increasing the QoI Dimension
The iterative method requires no such pre-processing of the data or the use of any specialized method for learning a density on a lower-dimensional manifold. In this case, the iterative algorithm terminated at the 43rd epoch due to the relative KL tolerance being reached. The results of the final epoch are shown in Figure 10. Since each epoch now includes three iterations, each iteration of the last epoch is shown in the top row of plots while the KDE of the data-generating distribution is omitted. However, comparing these to the prior top-row of plots involving two QoI (whether the iterative or joint results), we observe similar, if not better, qualitative agreement with the structure of the KDE of the data-generating distribution. This indicates that the inclusion of additional QoI that are “geometrically redundant” does not deteriorate the iterative DCI solution whereas the joint solution is not even computable unless more sophisticated methods are utilized as mentioned above.
@@ 6.2.1 Numerical Setup
where $\overline{Y}$ is the mean field and $\lambda_{i}$ are mutually uncorrelated random variables with zero mean and unit variance [21, 42]. We refer to $\lambda_{i}$ as the KL modes, which should not be confused with the KL divergence. The eigenvalues, $\eta_{i}$, and eigenfunctions, $f_{i}$, are computed numerically using the following covariance function,
$C_{Y}({\mathbf{x}},\overline{{\mathbf{x}}})=\sigma_{Y}^{2}\exp\left[-\frac{(x_{1}-\overline{x}_{1})^{2}}{2\eta_{1}}-\frac{(x_{2}-\overline{x}_{2})^{2}}{2\eta_{2}}\right],$
where $\sigma_{Y}$ and $\eta_{i}$ denote the variance and correlation length in the $i$th spatial direction, respectively. For a given correlation length, the expansion is typically truncated once a sufficient fraction of the energy in the eigenvalues is retained [46, 21]. However, we set the correlation length to $0.1$ in each spatial direction and truncate the expansion at 100 terms for the sake of demonstration.
To approximate solutions to the PDE in (21) we use the open source finite element software MrHyDE [45, 44] to implement a mixed finite element discretization with a physics-compatible discretization using the lowest-order Raviart-Thomas elements to approximate the velocity field, defined as $\mathbf{u}=-K\nabla p$, and a piecewise constant approximation of the pressure on a ($100\times 100$) spatial grid of $\Omega$. To define our QoI, we generate 1000 random sensor locations throughout $\Omega$ and measure both the pressure and velocity at these locations. Only a small subset of these were actually used for the numerical results in this paper.
@@
In the results shown in this work, we demonstrate the utility of the iterative DCI solution by considering two different GSIPs defined by a 12-dimensional QoI map that contains the two types of measurements (pressure and velocity) obtained from randomly placed sensors within the domain. Specifically, the first four sensors are assumed to measure both the pressure and the norm of the velocity field, which define eight of the QoI, and the last four sensors are assumed to measure only the pressure, which defines the other four QoI. The locations of these eight sensors are shown in Figure 11. In the supplementary material, we include the scripts to run MrHyDE and regenerate the full datasets. The datasets included in the supplementary material allow the user to define various QoI maps of different dimension from 1000 randomly generated sensor locations for which either pressure or velocity (or both) measurements can be chosen. The eight sensor locations chosen to construct the 12-dimensional QoI map in this example are simply the first eight randomly generated locations of this dataset. For the first GSIP considered, we iterate through all 12 components of this map. For the second GSIP considered, we assume that the pressure and velocity data are obtained simultaneously. In that case, the four separate 2-dimensional joint observed and predicted densities are estimated and utilized in the iterative algorithm defining a total of eight QoI subspaces that are iterated over in the algorithm. By utilizing the iterative DCI approach in both of these GSIPs, we avoid the complications inherent in approximating non-parametric distributions in high-dimensions with KDEs. As in the previous example, all densities are approximated with a standard Gaussian KDE with Scott’s rule used to determine the bandwidth parameter [39]. Since there are significantly more push-forward constraints to simultaneously satisfy in this example compared to the previous example, we relax the absolute KL divergence tolerance to ${1}\mathrm{E}{-6}$ in this example while leaving the remaining stopping criteria for the algorithm identical to those utilized in the previous example. Additionally, since the scales of the pressure and velocity data are significantly different, we apply a standard scaling to all QoI data before any KDEs are applied so that all visualizations of the QoI densities are at similar scales.

### v2

@@ 1 Introduction
A key objective of modern computational science and engineering is the rigorous quantification of uncertainty in physics-based models, driven by the desire to produce data-informed predictions. Central to this effort is the formulation and solution of stochastic inverse problems (SIPs), which seek to constrain input uncertainties using observational data. The SIP is broadly defined in the context of this work as constructing a probability measure on an input space ($\Lambda$) of a measurable quantities of interest (QoI) map ($\phi$) that is subject to a push-forward constraint associated with some specified observed probability measure ($P_{\text{obs}}$) on the output space ($\mathcal{D}$). More formally, the solution is a probability measure $P$ on $\Lambda$ such that $P\circ\phi^{-1}=P_{\text{obs}}$. In other words, the solution is a pullback measure of the observed distribution, assigning the correct probability to events in the parameter space based on observed frequencies in the data space. In that sense, the solution is naturally interpreted as being a data-generating distribution that is consistent with the observed output statistics and the assumed forward model.
While the standard SIP considers a single (possibly vector-valued) QoI map, many practical applications involve conditioning parameter uncertainties on multiple, distinct observational datasets associated with separate QoI maps. Such situations arise when different experimental campaigns lead to asynchronous data for which no joint structures are observed or known simultaneously across every QoI component from each QoI map. Additionally, in many realistic applications, even if joint structures were known across every QoI component, methods for estimating densities on such subspaces from finite data samples have practical dimension limitations that motivate a reformulation of the SIP on multiple, low-dimensional QoI subspaces. This leads to the Generalized Stochastic Inverse Problem (GSIP), which requires finding a single probability measure $P$ that simultaneously satisfies $k$ distinct push-forward constraints, $P\circ\phi_{i}^{-1}=P_{\text{obs},i}$, for $i=1,\dots,k$, where $\phi_{i}$ denotes the $i$th (possibly vector-valued) QoI map and $P_{\text{obs},i}$ is the associated observed probability measure. The GSIP presents unique challenges compared to the SIP, particularly because the full joint observed distribution across all output spaces ($\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{k}$ where $\mathcal{D}_{i}$ is the output space associated with $\phi_{i}$) is generally not assumed to be available or necessary for solution. Subsequently, solving the GSIP requires an approach that resolves these constraints individually and iteratively.
@@
To help situate the main contributions of this work within the broader landscape of inverse problems, especially those based on entropy-regularized methods, we note some conceptual connections to classical methods. Specifically, the iterative updates resemble adaptive/iterative importance sampling schemes, and the cycling over multiple marginal constraints is motivated by classical iterative proportional fitting. We also note that the convergence result identifying the limit as the I-projection also has a clear information-geometric interpretation: among all parameter measures that satisfy the push-forward constraints, the algorithm selects the one closest to the chosen reference measure $P_{\text{init}}$. When $P_{\text{init}}$ is uniform on the feasible set, this provides a maximum-entropy characterization under the imposed constraints.
@@ 2 Data-Consistent Inversion (DCI): Background, Notation, and Terminology
Let $P_{\text{init}}$ denote a probability measure on $(\Lambda,\mathcal{B}_{\Lambda})$ and $P_{\text{pred}}:=P_{\text{init}}\circ\phi^{-1}$ denote its push-forward measure on $(\mathcal{D},\mathcal{B}_{\mathcal{D}})$, then the disintegration of $P_{\text{init}}$ yields
@@ 3 DCI Solutions Minimize f-Divergences
It is worth noting that by varying $P_{\text{init}}$, we can generate a range of possible solutions, $P_{\text{up}}$, implying that any pullback measure can be recovered by a well chosen initial distribution. It may also be necessary to vary $P_{\text{init}}$ to ensure that $P_{\text{init}}\ll P_{\text{up}}$, which ensures that the space of potential solutions is nonempty. This flexibility in the choice of an initial distribution ultimately allows for a more comprehensive exploration of the solution space. Although varying $P_{\text{init}}$ is not the focus of this work, it will be the topic of a future work as mentioned in Section 7.
Once $P_{\text{init}}$ is chosen, we prove below that the DCI solution is optimal in the sense of minimizing the f-divergences over the space of all probability measures that are absolutely continuous with respect to $P_{\text{init}}$ and whose push-forward measure is equal to $P_{\text{obs}}$. The f-divergences [33] are commonly employed in probability and information theory to quantify the difference between probability measures often in the context of determining optimal parameters or hyper-parameters of density models, e.g., see [25, 5]. A key feature of f-divergences are their flexibility, as they generalize and connect many well-known divergence measures, such as the Kullback-Leibler (KL) divergence and total variation distance [27, 41, 5, 1]. Utilizing the notation of this work, the f-divergence is defined as:
@@
where $P\ll P_{\text{init}}$ and $f:[0,\infty)\to\mathbb{R}$ is a convex function satisfying $f(1)=0$ and $f(0)=\lim_{t\rightarrow 0^{+}}f(t)$. When $f(t)=t\log t$, we obtain the KL divergence.
@@
$P(\phi^{-1}(B))=P_{\text{obs}}(B)=\int_{B}\,dP_{\text{obs}}(\omega)=\int_{B}\cfrac{dP_{\text{obs}}}{dP_{\text{pred}}}(\omega)dP_{\text{pred}}(\omega).$
@@ 4 The Generalized Stochastic Inverse Problem
It is self-evident that if each $P_{\text{obs},i}$ is the induced push-forward measure of some probability measure defined on $(\Lambda,\mathcal{B}_{\Lambda})$, then a solution is guaranteed to exist. We make such an assumption in this work and refer to this measure as the data-generating distribution (denoted as $P_{\text{DG}}$), given its role in inducing the observed probability measures. Such an assumption is reasonable in practice when observed measures are associated with observed data, but such an assumption may be violated if observed measures are specified as target measures a priori of any data collection.
We briefly emphasize several important differences between the GSIP and the SIP. First, for the SIP solved by DCI, only a single (possibly vector-valued) map $\phi$ is considered. By contrast, there are multiple (possibly vector-valued) maps considered in the GSIP. Moreover, for the GSIP, it is possible that some components of a map $\phi_{i}$ may also appear in other maps $\phi_{j}$. This is related to a second key difference in terms of how uncertainties in output data are assumed to be quantified in the SIP and GSIP. In the SIP, a single joint observed measure is assumed to be defined on the full data space defined by all the outputs of the map $\phi$. However, the GSIP does not assume that the full joint observed distribution is available on $\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{k}$. A final important note is that a predictability assumption is sufficient to guarantee the DCI solution exists to the SIP (e.g., that $P_{\text{obs}}$ is absolutely continuous with respect to $P_{\text{pred}}=P_{\text{init}}\circ\phi^{-1}$). While an analogous predictability assumption is also required for the solution of the GSIP, a further assumption is needed to guarantee that each of the separate $k$ push-forward constraints of (14) are simultaneously satisfied, which is handled in this work by the additional assumption of a data-generating distribution.
@@ 5 An Iterative DCI Approach
Before we present the proof, we provide high-level comments to help provide some intuition and orient the reader on the steps involved. The proof we provide shows the $k=3$ case, and we note that the approach we provide is extended to the $k>3$ cases naturally. The first step in the proof constructs a product data space from the $k$ push-forward constraints along with a family of push-forward probability measures since each constraint defines a known marginal on this space. The second step in the proof involves expressing the Radon-Nikodym derivatives of the push-forward measures with respect to the push-forward of $P^{0}=P_{\text{init}}$, which provides a representation in the form of an IPFP on the product data space. To help build some intuition about the structure of these Radon-Nikodym derivatives and the relation to the IPFP method, suppose we equip $\mathbb{R}^{3}$ with the initial density $f^{0}(x,y,z)$ but that we wish to adjust this joint density so that the $x$-marginal is changed from $f^{0}_{1}(x):=\int_{\mathbb{R}^{2}}f^{0}(x,y,z)\,dy\,dz$ to $f_{1}(x)$. This is easily accomplished through the following formula (assuming the support of $f_{1}(x)$ is a subset of the support of $f^{0}_{1}(x)$)
@@ 6.1.1 Numerical Setup
For the algorithm, the stopping criteria were defined as follows. The maximum of number of epochs was set to 100, which was determined by trial-and-error to ensure that the algorithm stopped due to other criteria. The tolerance was set to 0.1 for the numerical diagnostic, i.e., if the sample average of the $r$-values deviates by more than 10% from the theoretically expected value of unity, then the algorithm terminates. This value of 0.1 was chosen based on the prior experience of the authors regarding typical variation seen in the diagnostic for DCI problems where the predictability assumption is satisfied. The tolerance for the KL divergence of all QoI marginals associated with the observed distribution and push-forward of the updated distribution after each epoch was set to ${1}\mathrm{E}{-7}$. In other words, this tolerance is reached at epoch $k$ only if every QoI marginal associated with the observed distribution and the push-forward of the updated distribution is less than ${1}\mathrm{E}{-7}$. This value was chosen based on some numerical experiments so that it was smaller than typically observed KL divergence values between KDE estimates obtained from different sample sets taken from the same distribution. Finally, the relative tolerance for the difference in such KL divergences obtained between successive epochs was set to ${1}\mathrm{E}{-2}$. In other words, this tolerance is reached at epoch $k+1$ only if the KL divergences of every QoI marginal associated with the observed and updated distributions are within $1\%$ of those computed at epoch $k$.
@@ 6.1.2 Iterative Results
The relative tolerance stopping criteria was triggered at the 46th epoch where the push-forward of the updated distribution at this epoch produced marginal QoI densities that both had KL divergences from the associated observed marginals that were within 1% from the values computed from the 45th epoch. We illustrate some of the results in Figures 3-5. For reference, the top-right plot in each of these figures shows the KDE of the data-generating samples shown as the orange pluses in the left plot of Figure 2. The top-left and top-middle plots in each of these figures show the weighted KDE computed on the initial samples shown as the blue circles in Figure 2 with weights determined at the first (left) and second (middle) iteration of each epoch. The bottom two plots in each of these figures show the KDE estimates of the marginal densities on $Q_{1}$ (left) and $Q_{2}$ (right) associated with the observed distribution and push-forward of the updated distribution at the end of each epoch.
@@
Numerical evidence of convergence is further supported by Figure 6, which shows the rapid decrease in the KL divergence between the QoI marginals associated with the observed distribution and the push-forward of the updated distribution at each iteration. Here, we show results for each iteration (of which there are 92 total) instead of each epoch (of which there are 46) to highlight the impact of alternating between the $Q_{1}$ and $Q_{2}$ distributions at each iteration within an epoch that produce oscillations in the decay trend. The magnitude of the oscillations is observed to quickly decay before essentially becoming flat with minor oscillations smaller than ${1}\mathrm{E}{-6}$ in magnitude after 30-40 iterations (i.e., after 15-20 epochs).
@@ 6.1.4 Increasing the QoI Dimension
The iterative method requires no such pre-processing of the data or the use of any specialized method for learning a density on a lower-dimensional manifold. In this case, the iterative algorithm terminated at the 43rd epoch due to the relative KL tolerance being reached. The results of the final epoch are shown in Figure 10. Since each epoch now includes three iterations, each iteration of the last epoch is shown in the top row of plots while the KDE of the data-generating distribution is omitted. However, comparing these to the prior top-row of plots involving two QoI (whether the iterative or joint results), we observe similar, if not better, qualitative agreement with the structure of the KDE of the data-generating distribution. This indicates that the inclusion of additional QoI that are “geometrically redundant” does not deteriorate the iterative DCI solution in this example whereas the joint solution is not even computable unless more sophisticated methods are utilized as mentioned above.
@@ 6.2.1 Numerical Setup
where $\overline{Y}$ is the mean field and $\lambda_{i}$ are mutually uncorrelated random variables with zero mean and unit variance [21, 42]. Various QoI for this problem are considered from the evaluation of the pressure and/or norms of the velocity field, given by $K\nabla p$, at various points within the physical domain $\Omega$ while the parameters are given by the $\lambda$ inputs to the permeability field. We refer to $\lambda_{i}$ as the KL modes, which should not be confused with the KL divergence. The eigenvalues, $\eta_{i}$, and eigenfunctions, $f_{i}$, are computed numerically using the following covariance function,
$C_{Y}({\mathbf{x}},{\mathbf{x}}^{\prime})=\sigma_{Y}^{2}\exp\left[-\frac{(x_{1}-{x}^{\prime}_{1})^{2}}{2\zeta_{1}}-\frac{(x_{2}-{x}^{\prime}_{2})^{2}}{2\zeta_{2}}\right],$
where $\sigma_{Y}$ and $\zeta_{i}$ denote the variance and correlation length in the $i$th spatial direction, respectively. For a given correlation length, the expansion is typically truncated once a sufficient fraction of the energy in the eigenvalues is retained [46, 21]. However, we set the correlation length to $0.1$ in each spatial direction and truncate the expansion at 100 terms for the sake of demonstration since these were observed to produce significant variability in both the permeability fields and ranges of potential QoI values.
To approximate solutions to the PDE in (21) we use the open source finite element software MrHyDE [45, 44] to implement a mixed finite element discretization with a physics-compatible discretization using the lowest-order Raviart-Thomas elements to approximate the velocity field, defined as $\mathbf{u}=-K\nabla p$, and a piecewise constant approximation of the pressure on a ($100\times 100$) spatial grid of $\Omega$. In Figure 11, we plot two random realizations of the permeability field generated using the truncated KL expansion.
Figure 11: Two random realizations of the permeability field generated using the truncated KL expansion.
Correspondingly, in Figures 12 and  13 we plot the pressure fields and the magnitude of the velocity fields for each of these random realizations.
Figure 12: The approximated pressure fields corresponding to the two random realizations of the permeability field generated using the truncated KL expansion.
Figure 13: The magnitude of the approximated velocity fields corresponding to the two random realizations of the permeability field generated using the truncated KL expansion.
To define our QoI, we generate 1000 random sensor locations throughout $\Omega$ and measure both the pressure and velocity at these locations. Only a small subset of these were actually used for the numerical results in this paper.
@@
In the results shown in this work, we demonstrate both the utility and flexibility of the iterative DCI solution by considering two different GSIPs defined by a 12-dimensional QoI map that contains the two types of measurements (pressure and velocity) obtained from randomly placed sensors within the domain. Specifically, the first four sensors are assumed to measure both the pressure and the norm of the velocity field, which define eight of the QoI, and the last four sensors are assumed to measure only the pressure, which defines the other four QoI. The locations of these eight sensors are shown in Figure 14. In the supplementary material, we include the scripts to run MrHyDE and regenerate the full datasets. The datasets included in the supplementary material allow the user to define various QoI maps of different dimension from 1000 randomly generated sensor locations for which either pressure or velocity (or both) measurements can be chosen. The eight sensor locations chosen to construct the 12-dimensional QoI map in this example are simply the first eight randomly generated locations of this dataset. For the first GSIP considered, we iterate through all 12 components of this map. For the second GSIP considered, we assume that the pressure and velocity data are obtained simultaneously. In that case, the four separate 2-dimensional joint observed and predicted densities are estimated and utilized in the iterative algorithm defining a total of eight QoI subspaces that are iterated over in the algorithm. By utilizing the iterative DCI approach in both of these GSIPs, we avoid the complications inherent in approximating non-parametric distributions in high-dimensions with KDEs. As in the previous example, all densities are approximated with a standard Gaussian KDE with Scott’s rule used to determine the bandwidth parameter [39]. Since there are significantly more push-forward constraints to simultaneously satisfy in this example compared to the previous example, we relax the absolute KL divergence tolerance to ${1}\mathrm{E}{-6}$ in this example while leaving the remaining stopping criteria for the algorithm identical to those utilized in the previous example. Additionally, since the scales of the pressure and velocity data are significantly different, we apply a standard scaling to all QoI data before any KDEs are applied so that all visualizations of the QoI densities are at similar scales.